OpenStudy (anonymous):

Part 1: Explain, in complete sentences, how you would use the substitution method to solve the following system of equations. (3 points) Part 2: Provide the solution to the system. (1 point) 5x – y = –1 7x – 3y = 13

OpenStudy (earthcitizen):

first i would make y the subject in eqn 1 and plug it into eqn 2 to find x. Then use x to find y

OpenStudy (earthcitizen):

y=5x+1...eqn 1 7x-3(5x+1) =13, 7x-15x-3=13 => x=-2

OpenStudy (earthcitizen):

\[\therefore y=5(-2)+1=-9\]

OpenStudy (anonymous):

this is from virtual school isn't it

OpenStudy (earthcitizen):

virtual school?

OpenStudy (anonymous):

yeah I'm taking algebra 1 online and its the exact same question and format that i had on the test

OpenStudy (earthcitizen):

how was the test ? did you make it ?

OpenStudy (anonymous):

no i didn't make it i took it though

OpenStudy (earthcitizen):

yh, did you pass it ?

OpenStudy (anonymous):

yeah... i think why does it matter

OpenStudy (earthcitizen):

never mind.Good job tho

OpenStudy (anonymous):

ok... i took it a while ago